# Semi-Supervised In-Context Learning: A Baseline Study

### v1

@@ Abstract
Most existing work in data selection for In-Context Learning (ICL) has focused on constructing demonstrations from ground truth annotations, with limited attention given to selecting reliable self-generated annotations. In this work, we propose a three-step semi-supervised ICL framework: annotation generation, demonstration selection, and semi-supervised inference. Our baseline, Naive-SemiICL, which prompts select high-confidence self-generated demonstrations for ICL prompting, outperforms a 16-shot baseline by an average of 9.94% across 16 datasets. We further introduce IterPSD, an annotation approach that refines pseudo-demonstrations iteratively, achieving up to 6.8% additional gains in classification tasks. Lastly, we reveal a scaling law for semi-supervised ICL, where models achieve optimal performance with over 1,000 demonstrations. Code is available at: https://anonymous.4open.science/r/semi-supervised-icl-FA07
@@
Semi-Supervised In-Context Learning: A Baseline Study
@@ 1 Introduction
In-Context Learning (ICL) has emerged as a powerful paradigm in natural language processing, enabling Language Models (LMs) to learn, adapt, and generalize from examples presented within their input context. This approach eliminates the need for extensive retraining and parameter modifications, facilitating more flexible and efficient learning Brown et al. (2020); Min et al. (2022); Agarwal et al. (2024); Fang et al. (2025). The effectiveness of ICL is contingent on the quality of the annotated examples, which encompasses factors such as diversity He et al. (2023), representativeness Yang et al. (2023), and ordering Lu et al. (2021). While significant efforts have been dedicated to selecting high-fidelity demonstrations using ground-truth annotations, the selection of reliable self-generated annotations for ICL remains underexplored. Existing methods are often constrained to specific tasks Zhang et al. (2023); Mamooler et al. (2024); Li et al. (2024a) and rely on task-specific heuristics Zhang et al. (2023). For instance, Auto-CoT Zhang et al. (2023) enhances demonstration quality by ensuring diversity and applying heuristic-based filtering rules, such as constraints on the size of reasoning steps. PICLe Mamooler et al. (2024) leverages LLM self-verification to discard demonstrations with incorrect named entity types, while SAIL Li et al. (2024a) filters out predictions that fail to translate back into the original input.
@@
To advance domain-agnostic approaches for self-annotated demonstration selection, we propose a three-step framework (Figure 1) comprising: (1) annotation generation, (2) demonstration selection, and (3) semi-supervised inference. Drawing inspiration from semi-supervised learning, we term this framework Semi-Supervised ICL. We introduce a simple baseline, Naive-SemiICL, which selects high-confidence self-generated annotations as demonstration examples. Further extending this analogy, we propose IterPSD, an iterative approach for generating pseudo-demonstrations that progressively refines prompts by incorporating self-generated annotations at each step.
We systematically evaluate Naive-SemiICL against a robust 16-shot baseline using three confidence scoring methods across 16 datasets spanning 9 tasks. Experimental results indicate that Naive-SemiICL consistently outperforms the 16-shot baseline, achieving an average performance improvement of 9.94% on GPT-4o-mini (Figure 2). Furthermore, we demonstrate that IterPSD enhances semi-supervised ICL performance in five classification tasks, yielding gains of up to 6.8% (Table 1). Additionally, we identify a scaling law in semi-supervised ICL, where inference accuracy improves with increasing pseudo-demonstration examples, reaching optimal performance beyond 1,000 examples (Figure 3).
Figure 2: Comparison of GPT-4o-mini (top) and GPT-4o (bottom) performance across multiple datasets using three different methods: Few-Shot, Naive-SemiICL (Naive), and Naive-SemiICL without data selection (Unfiltered).
@@ 2 Method
In this section, we establish the framework of Semi-Supervised ICL, which consists of three phases: ① pseudo-demonstration generation, ② demonstration selection and ③ semi-supervised inference. Then, we propose a simple baseline for semi-supervised ICL, Naive-SemiICL, that generates pseudo-demonstrations in one iteration and filters out examples with a low confidence score. Building on top of Naive-SemiICL, we propose an iterative method for generating pseudo-demonstrations, IterPSD, that progressively enhances its prompt by incorporating self-generated annotations during the demonstration generation process.
@@ 2.1 Semi-Supervised ICL
Semi-Supervised ICL. Beyond ground-truth annotations, Semi-Supervised ICL leverages un-annotated data to enrich demonstrations. The setting assumes the availability of a ground-truth dataset $\mathcal{D}_{g}=\{(x_{i},y_{i})\}^{N_{l}}$, usually small in quantity, alongside a large pool of un-annotated data $\mathcal{X}_{u}={x_{i}}^{N_{u}}$. Semi-Supervised ICL augments the limited pool of ground-truth examples by generating pseudo-demonstrations $\mathcal{D}_{\mathrm{PSD}}$ from the un-annotated data. Formally,
@@
where $(\tilde{y},\tilde{r},\tilde{c})=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E}_{g},x)$ are generated by the LLM in an ICL fashion using a set of ground-truth demonstrated ns $\mathcal{E}_{g}\subseteq\mathcal{D}_{g}$. Low-confidence examples are then filtered out according to a confidence threshold $\lambda$:
@@
Figure 3: Pseudo-demonstration scaling (PSD) performance of Naive-SemiICL on GPT-4o (top) and GPT-4o-mini (bottom) across multiple datasets. The blue solid line represents Naive-SemiICL performance, and the dashed gray line represents the few-shot baseline.
@@ 2.2 A Simple Semi-Supervised ICL Baseline
We proposed a simple method that generates the pseudo-demonstrations in one iteration (Algorithm 1). We dub the method, together with the rest of the Semi-Supervised ICL framework, Naive-SemiICL. The method simply goes through the un-annotated data and generates a prediction, a rationale and a confidence score for each input. As the simplest form of Semi-Supervised ICL, it provides effective in-context learning signals by filtering out low-quality pseudo-demonstrations. We experiment with three commonly used confidence measures on 15 datasets spanning 9 tasks and show that this simple baseline consistently outperforms a strong 16-shot ICL baseline (Section 4.1).
@@ 2.3 Iterative Pseudo-Demonstration Generation
Encouraged by the success of Naive-SemiICL, we explore whether pseudo-demonstrations can enhance the accuracy of subsequent pseudo-demonstration generation. We propose IterPSD (Algorithm 2), an iterative method for generating pseudo-demonstrations that
@@
    recursively adds newly generated pseudo-demonstrations to its own prompt before reaching maximum number of allowed demonstrations (Line 6), and
@@
    re-sample the most confident pseudo-demonstrations according to a confidence threshold $\lambda$ from all previously annotated instances once the size of the demonstration reaches a limit (Line 12).
In each iteration, IterPSD samples and annotates $K$ un-annotated examples before applying a filtering step. The generated pseudo-demonstrations are recursively accumulated and fed back into the LLM to generate additional pseudo-demonstrations (Line 10). To mitigate performance degradation caused by long context lengths, we impose an upper limit $\kappa$ on the number of self-fed pseudo-demonstrations. Once this limit is reached, the $\kappa$ most confident pseudo-demonstrations are resampled from the entire generated set, ensuring that only high-quality examples are retained (Line 7).
@@
To enhance the accuracy of pseudo-demonstration annotations, we introduce the $\epsilon$-Random Sampler (Algorithm 3), a sampling strategy that selects both similar and diverse examples from the un-annotated pool. At each iteration, a proportion $(1-\epsilon)$ of examples is chosen based on their cosine similarity to the nearest previously annotated instances (Line 6) [^2] We compute cosine similarity using vector embeddings generated by OpenAI’s text-embedding-3-large., ensuring that each selected example is closely related to an existing annotation. The underlying intuition is that the model achieves higher annotation accuracy on similar examples, thereby improving the quality of subsequent annotations. This approach aligns with curriculum learning Soviany et al. (2021), a widely used technique in semi-supervised learning that prioritizes simpler, well-understood examples before progressively annotating more challenging ones.
@@
Mitigating Confirmation Bias.
To maintain annotation quality, we ensure that at least half of the data ($\epsilon\geq 0.5$) is sampled randomly (Line 7) [^3] We found best performing $\epsilon$ to be $0.8$ in our experiments.. Without this, annotation accuracy degrades significantly. When $\epsilon=0$, selections are exclusively based on similarity to previously annotated examples. Pseudo-demonstrations become overly homogeneous, leading to bias in ICL predictions. This phenomenon closely parallels confirmation bias in semi-supervised learning Arazo et al. (2019); Zou and Caragea (2023), highlighting a strong connection between Semi-Supervised ICL and traditional semi-supervised learning frameworks.
@@ 3 Experimetnal Setup
  • Classification. We include BANKING77 Casanueva et al. (2020) and CLINC Larson et al. (2019) for intent classification, CLINC(D) for domain classification, FewEvent Deng et al. (2020) for few-shot event classification, and FP Malo et al. (2013) for sentiment analysis.
  • Translation. We evaluate Naive-SemiICL’s ability to translate English into low-resource language using the FLORES200 dataset Costa-Jussà et al. (2022). In particular, we evaluate on six languages: Bemba, Fijian, Faroese, Tuvan, Venetian, and Sardinian.
  • Reasoning. We evaluate our method’s reasoning skills on GPQA Rein et al. (2024), a graduate-level scientific reasoning task, LiveBench Math White et al. (2025), a competition-level mathematics problems dataset, and three tasks from BigBenchHard Suzgun et al. (2022): Logical7, which tests LLM’s spatial reasoning ability, Geometric Shapes, which tests ability to interpret and identify geometric figures, and Date, which evaluates time manipulation skills.
We detail how we split these datasets for demonstration and evaluation in Appendix B.1.
Evaluation Metrics. For all classification and reasoning tasks, we report accuracy as the performance metric. We evaluate the equivalence of latex-style math outputs on LiveBench Math using the parser reported in Gao et al. (2024). For translation tasks, we report ChrF++ score Popović (2015) in its default configuration as implemented by TorchMetricsDetlefsen et al. (2022) following Agarwal et al. (2024). We report the average and standard error of three trials OpenAI (2024) on baseline results (Section 4.1). The rest of the results are collected over one trial.
Baselines. On classification and translation tasks, we compare Naive-SemiICL to a $16$-shot baseline. We report the best performing $k$-shot baseline where $k\leq 16$ on reasoning tasks, as the best-performing baseline does not necessarily take 16 demonstrations (see details in Appendix B.3). During pseudo-demonstration generation, we use the same ground truth demonstration examples that the baseline uses. We generate the label/translation directly on classification and translation tasks and leverage Chain-of-Thought prompting to generate rationale on reasoning tasks (see Appendix A for the CoT prompt). We experiment with different pseudo-demonstration sizes $k_{u}$ for Naive-SemiICL: $k_{u}\in\{32,64,100,150,200\}$ on classification tasks, and $k_{u}\in\{32,64,100,150,200\}$ on translation tasks.
@@
We primarily evaluate three confidence metrics: Verbalized Confidence, Entropy, and Self-Consistency, as defined below.
Verbalized Confidence. Verbalized Confidence Xiong et al. (2024) prompts the LLM to generate the confidence score as part of its response. See Appendix A for how we induce the Verbalized Confidence scores from LLMs in prompts.
Entropy. Entropy Shannon (1948) estimates the uncertainty of generated content using the token probability $P(w_{i}\mid w_{<i})$, where $w_{i}$ denotes the generated tokens and $w_{<i}$ represents the preceding tokens in the prompt:
@@
We find Entropy unsuitable for estimating uncertainty on classification tasks, as it predominantly returns a confidence score of one, making the data selection step redundant.

### v2

# Scaling Laws for Many-Shot In-Context Learning with Self-Generated Annotations

@@ Abstract
The high cost of obtaining high-quality annotated data for in-context learning (ICL) has motivated the development of methods that use self-generated annotations in place of ground-truth labels. While these approaches have shown promising results in few-shot settings, they generally do not scale to many-shot scenarios. In this work, we study ICL with self-generated examples using a framework analogous to traditional semi-supervised learning, consisting of annotation generation, demonstration selection, and in-context inference. Within this framework, we propose a simple baseline that outperforms ground-truth ICL in zero-shot, few-shot, and many-shot settings. Notably, we observe a scaling law with this baseline, where optimal performance is achieved with more than 1,000 demonstrations. To fully exploit the many-shot capabilities of semi-supervised ICL, we introduce IterPSD, an iterative annotation approach that integrates iterative refinement and curriculum pseudo-labeling techniques from semi-supervised learning, yielding up to 6.8% additional gains on classification tasks. Code is available at: https://anonymous.4open.science/r/semi-supervised-icl-FA07
@@
Scaling Laws for Many-Shot In-Context Learning with Self-Generated Annotations
@@ 1 Introduction
In-context learning (ICL) has emerged as a powerful paradigm in natural language processing, enabling language models (LMs) to learn, adapt, and generalize from examples presented within their input context. This approach eliminates the need for extensive retraining and parameter modifications, facilitating more flexible and efficient learning Brown et al. (2020); Min et al. (2022); Agarwal et al. (2024); Fang et al. (2025). The high cost of obtaining high-quality annotated data for ICL has motivated the development of methods Zhang et al. (2023); Li and Qiu (2023); Mamooler et al. (2024); Li et al. (2024a); Chen et al. (2023) that use self-generated annotations in place of ground-truth labels. However, previous research has not examined ICL performance with self-generated annotations in many-shot settings. Recently, Agarwal et al. (2024) established a scaling law, showing that ICL performance improves with the number of demonstrations—up to thousands of examples. Inspired by this finding, we pose the following question:
Research Question: Can we scale ICL performance using self-generated demonstrations up to thousands of examples as well?
We systematically investigate this question under a three-step framework (Figure 1): ① annotation generation, ② demonstration selection, and ③ semi-supervised inference, which we term Semi-Supervised ICL. We first introduce a simple baseline, Naive-SemiICL, which annotates unlabeled data in a single iteration, scoring each annotation using the LLM’s verbalized confidence. Naive-SemiICL consistently outperforms ICL baselines in zero-shot, few-shot, and many-shot settings, as well as prior methods. We highlight that Naive-SemiICL achieves optimal performance with 1000 demonstrations on certain tasks (Figure 2).
With potentially thousands of self-annotated examples in the prompt, each demonstration can be viewed as a dataset, which motivates the following question:
Research Question: In what ways can techniques from traditional semi-supervised learning be leveraged to improve ICL performance?
We address this question by proposing IterPSD, an iterative approach that progressively refines pseudo-demonstration quality by incorporating self-generated annotations at each iteration. IterPSD further improves semi-supervised ICL performance on five classification tasks, achieving gains of up to 6.8% (Table 3).
Figure 2: Scaling trend of Naive-SemiICL on classification and translation tasks with GPT-4o and GPT-4o-mini. The dashed gray line represents the few-shot baseline. Both model exhibits a scaling trend on most tasks. All experiments are performed with a ground truth budget of $k_{l}=16$.
@@ 2 Method
In this section, we establish the framework of Semi-Supervised ICL, which consists of three phases: ① pseudo-demonstration generation, ② demonstration selection, and ③ semi-supervised inference. We then propose a simple baseline for Semi-Supervised ICL, Naive-SemiICL, which generates pseudo-demonstrations in a single iteration and filters out examples with low confidence scores. Building on Naive-SemiICL, we introduce an iterative method, IterPSD, that progressively improves the prompt by incorporating self-generated annotations during the demonstration generation process.
@@ 2.1 Semi-Supervised ICL
Semi-Supervised ICL. Beyond ground-truth annotations, Semi-Supervised ICL leverages unannotated data to enrich demonstrations. The setting assumes the availability of a ground-truth dataset $\mathcal{D}_{g}=\{(x_{i},y_{i})\}^{N_{l}}$, usually small in quantity, alongside a large pool of unannotated data $\mathcal{X}_{u}={x_{i}}^{N_{u}}$. Semi-Supervised ICL augments the limited pool of ground-truth examples by generating pseudo-demonstrations $\mathcal{D}_{\mathrm{PSD}}$ from the un-annotated data. Formally,
@@
where $(\tilde{y},\tilde{r},\tilde{c})=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E}_{g},x)$ are generated by the LLM in an ICL fashion using a set of ground-truth demonstrated ns $\mathcal{E}_{g}\subseteq\mathcal{D}_{g}$. Low-confidence examples are filtered out according to a confidence threshold $\lambda$:
@@ 2.2 A Simple Semi-Supervised ICL Baseline
We propose a simple method that generates pseudo-demonstrations in a single iteration (Algorithm 1). We dub this method, along with the rest of the Semi-Supervised ICL framework, Naive-SemiICL. The method simply iterates over the unlabeled data for one iteration and generates a prediction, a rationale, and a confidence score for each input. As the simplest form of Semi-Supervised ICL, it provides effective in-context learning signals by filtering out low-quality pseudo-demonstrations. We experiment with three commonly used confidence measures on 16 datasets spanning 9 tasks, and show that this simple baseline consistently outperforms a strong 16-shot ICL baseline (Section 4.1).
@@ 2.3 Iterative Pseudo-Demonstration Generation
Encouraged by the success of Naive-SemiICL, we explore whether pseudo-demonstrations can enhance the accuracy of subsequent pseudo-demonstration generation. We propose IterPSD (Algorithm 2), an iterative method for generating pseudo-demonstrations that:
@@
    recursively adds newly generated pseudo-demonstrations to its own prompt until reaching the maximum number of allowed demonstrations (Line 6), and
@@
    re-samples the most confident pseudo-demonstrations according to a confidence threshold $\lambda$ from all previously annotated instances once the demonstration size reaches its limit (Line 12).
In each iteration, IterPSD samples and annotates $K$ unlabeled examples before applying a filtering step. The generated pseudo-demonstrations are recursively accumulated and fed back into the LLM to generate additional pseudo-demonstrations (Line 10). To mitigate performance degradation caused by long context lengths, we impose an upper limit $\kappa$ on the number of self-fed pseudo-demonstrations. Once this limit is reached, we resample the $\kappa$ most confident pseudo-demonstrations from the generated pseudo-demonstrations, ensuring that only high-quality examples are retained (Line 7).
Figure 3: Comparison of GPT-4o-mini (top) and GPT-4o (bottom) performance across multiple datasets using three different methods: Few-Shot, Naive-SemiICL (Naive), and Naive-SemiICL without filtering (Unfiltered).
@@
To enhance the accuracy of pseudo-demonstration annotations, we introduce the $\epsilon$-Random Sampler (Algorithm 3), a sampling strategy that selects both similar and diverse examples from the unannotated pool. At each iteration, a proportion $(1-\epsilon)$ of examples is chosen based on their cosine similarity to the nearest previously annotated instances (Line 6) [^2] We compute cosine similarity using vector embeddings generated by OpenAI’s text-embedding-3-large., ensuring that each selected example is similar to an existing annotation. The rest of the examples are chosen diversely in a clustering fashion similar to the one in Zhang et al. (2023). This approach that considers both the similarity and diversity aligns with curriculum learning Soviany et al. (2021) in semi-supervised learning, which facilitate self-training by balancing a mixture of confident and uncertain predictions.
Mitigating Confirmation Bias.
To maintain annotation quality, we find that at least half of the data ($\epsilon\geq 0.5$) should be sampled diversely (Line 7) [^3] We found the best performing $\epsilon$ to be $0.8$ in most of our experiments.. When $\epsilon=0$, selections are exclusively based on similarity to previously annotated examples. The Pseudo-demonstrations become homogeneous, leading to bias in ICL predictions. This phenomenon closely parallels confirmation bias in semi-supervised learning Arazo et al. (2019); Zou and Caragea (2023), highlighting a strong connection between Semi-Supervised ICL and traditional semi-supervised learning frameworks.
@@ 3 Experimetnal Setup
  • Classification. We include BANKING77 Casanueva et al. (2020), CLINC Larson et al. (2019), FewEvent Deng et al. (2020), and FP Malo et al. (2013).
  • Translation. We evaluate Naive-SemiICL’s ability to translate English into low-resource languages using 6 datasets from FLORES200 Costa-Jussà et al. (2022): Bemba, Fijian, Faroese, Tuvan, Venetian, and Sardinian.
  • Reasoning. We include 5 benchmarks spanning scientific, mathematical, and logical reasoning: GPQA Rein et al. (2024), LiveBench Math White et al. (2025), and three tasks from BigBenchHard Suzgun et al. (2022): Logical7, Geometric Shapes, and Date.
We describe these datasets in detail in Appendix B.3 and explain how we split the training and testing data in Appendix B.1.
Evaluation Metrics. For all classification and reasoning tasks, we report accuracy as the performance metric. We evaluate the equivalence of LaTeX-style mathematical outputs on LiveBench Math using the parser described in Gao et al. (2024). For translation tasks, we report the ChrF++ score Popović (2015) using its default configuration, as implemented in TorchMetrics Detlefsen et al. (2022), following Agarwal et al. (2024). We report the mean and standard error over three trials for baseline results OpenAI (2024) (Section 4.1). The remaining results are based on a single trial.
[…]
  • $k$-Shot ICL. The LLM is prompted with $k$ ground truth annotated examples, where $k$ ranges from 00 to $500$. Base on the number of ground truth annotation used, we divide our experiments into zero-shot (Table 1)), few-shot (Fig. 3), and many-shot (Fig. 6) settings.
  • Unfiltered SemiICL. To highlight the importance of confidence-based data selection, we include an unfiltered variant of Naive-SemiICL, which samples pseudo-annotations without applying the filtering step.
  • MoT. Li and Qiu (2023) We include MoT as a domain-specific baseline for reasoning tasks. Unlike MoT, Naive-SemiICL uses a simple one-step filtering mechanism for demonstration selection, whereas MoT requires querying the LLM for each example. Configuration details are provided in Appendix B.4.
  • Reinforced ICL. Agarwal et al. (2024) demonstrate that prompting the LLM with self-generated reasoning chains filtered by ground-truth answers can significantly improve ICL performance. This method serves as an upper bound on semi-supervised ICL performance on the reasoning tasks when the filtering mechanism is assumed to be perfect.
During our preliminary experiments, we found Auto-CoT to be uncompetitive on our reasoning datasets, as it relies on simple heuristics for data selection that are no longer effective. Since MoT includes all of Auto-CoT’s steps except its entropy- and semantic-based filters, we opted not to include Auto-CoT in our experiments.
@@
We primarily evaluate three confidence metrics: Verbalized Confidence which prompts the LLM to generate the confidence score (Table 7), Entropy, and Self-Consistency. Self-Consistency measures the confidence as the frequency of the most frequent answer, and entropy is defined as
@@
Task Zero-shot Naive Improv.
Classification Banking 61.50 78.00 26.8%
FewEvent 56.00 65.00 16.07%
CLINC 83.50 88.50 5.99%
CLINCD 59.50 61.00 2.52%
FP 91.00 94.50 3.85%
Translation Bemba 0.2437 0.2591 12.37%
Fijian 33.63 35.16 4.55%
Faroese 42.45 42.90 1.06%
Venetian 42.03 42.82 1.88%
Tuvan 16.17 18.40 13.79%
Sardinian 37.06 38.46 3.78%
Reasoning GPQA 36.36 38.38 5.55%
Math 35.48 36.58 3.10%
Logical7 65.00 72.00 10.77%
Shapes 56.00 60.00 7.14%
Date 40.00 65.00 62.5%
Table 1: Performance comparison of Zero-shot ICL and Naive-SemiICL. All experiments are done on GPT-4o-mini. For Naive-SemiICL, we report the best performing number of pseudo-demonstrations.
Hyperparameters.
Unless stated otherwise, we filter all generated pseudo-demonstrations using the confidence threshold at the 90th percentile. We discuss the hyperparameters of IterPSD in Appendix B.8.
Models.
We experiment with GPT-4o-mini and GPT-4o, checkpointed on 2024-07-18 and 2024-11-20, respectively, for all of our experiments. We discuss the computational cost associated with our experiments in Appendix B.2.
4 Empirical Analyses
4.1 Naive-SemiICL Consistently Beats Baselines
We first compare the performance of Naive-SemiICL with Verbalized Confidence to the few-shot baseline. For Naive-SemiICL, we report performance using the optimal number of pseudo-demonstrations $k_{u}$ for each task. The best-performing $k_{u}$ values are shown in Tables 5 and 6. Naive-SemiICL outperforms few-shot ICL on all tasks except CLINC(D), where it matches the baseline. Unfiltered SemiICL fails to match baseline performance in 20 out of 32 settings, highlighting the importance of the filtering step. A detailed breakdown of the best performance across different confidence scores is provided in Appendix C.
Method GPQA Math Logical7 Shapes Date
Naive-SemiICL 42.42 40.78 90.00 78.00 79.00
MoT 44.44 25.86 88.00 64.00 58.00
Reinforced ICL 54.54 42.63 93.00 78.00 89.00
Table 2: Comparison of Naive-SemiICL (Naive) and MoT on reasoning datasets using GPT-4o-mini.
We highlight the effectiveness of Naive-SemiICL in extremely low-resource settings through a zero-shot experimental design. We generate pseudo-demonstrations with no initial ground-truth demonstrations and compare Naive-SemiICL to zero-shot prompting. The performance gap between Naive-SemiICL and the zero-shot baseline depends solely on the quality of the filtering mechanism. As shown in Table 1, Naive-SemiICL outperforms the zero-shot baseline on all tasks in the benchmark, attaining an average improvement of 11.36% under GPT-4o-mini. This exceeds the average improvement of 9.94% in the 16-shot setting (Figure 3), suggesting that Naive-SemiICL is more effective resource-constrained conditions.
Additionally, we found Naive-SemiICL to be effective in high-resource settings. Figure 6 compares the performance of Naive-SemiICL and ground-truth ICL when $k_{l}\in\{64,100,500\}$ ground-truth examples are available. Across three tasks, Naive-SemiICL consistently outperforms the corresponding $k$-shot baselines. We observe diminishing returns in performance gains as the number of annotated demonstrations increases. On average, $k_{g}=64$ improves performance by 10.49% over the baseline, whereas $k_{g}=500$ yields only a 4.73% improvement across the three tasks. Combining these results, Naive-SemiICL is most effective when ground-truth data is scarce, although it can still be effective in high-resource settings.
On reasoning datasets, Naive-SemiICL outperforms MoT on all tasks except GPQA, as shown in Table 2. Surprisingly, the performance gap between the two methods is substantial on LiveBench Math, Shapes, and Date. We attribute this to two key differences between Naive-SemiICL and MoT: (1) MoT uses Entropy to filter low-quality demonstrations, which we show to be less reliable than Verbalized Confidence (see Table 4); and (2) in preliminary experiments, we found similarity-based retrieval to be less effective than diverse sampling. Naive-SemiICL samples diversely from a large pool of pseudo-demonstrations, which MoT is unable to do due to its requirement to query the LLM for each demonstration retrieval.
4.2 Scaling Law for Semi-Supervised ICL
We observe a scaling law for Semi-Supervised ICL, similar to the one reported in many-shot ICL Agarwal et al. (2024), on classification and translation tasks. We illustrate this trend in Figure 2. Across all configurations, Naive-SemiICL performance improves with larger demonstration sizes, although the point of peak performance varies. Both GPT-4o and GPT-4o-mini scale effectively across most tasks, typically peaking between 500 and 1,000 examples for classification tasks and between 100 and 200 examples for translation tasks. GPT-4o exhibits a more stable scaling trend than GPT-4o-mini on translation tasks, with performance peaking later and declining more gradually. With more available ground-truth data, we also observe scaling trends on BANKING77 and FewEvent (Figure 6).